can you please help solve 4c/8x*4x/22c
Found 2 solutions by vleith, stanbon:
Answer by vleith(2983)
Answer by stanbon(75887)

please help solve (4c/8x)*(4x/22c)
---
= (c/2x)*(2x/11c)
----
= c/(11c)
----
= 1/11
---------------
Cheers,
Stan H.
===============